Solving a linear system with substitution: -12x+y=15, 3x+2y=3
Answer by fractalier(6550)

From
-12x+y=15 and
3x+2y=3
let us solve the first one for y and substitute it in...
y = 12x + 15 and then
3x + 2(12x + 15) = 3
3x + 24x + 30 = 3
27x = -27
x = -1
and substituting back we get
3(-1)+2y=3
-3 + 2y = 3
2y = 6
y = 3
The solution is (-1, 3).